Solve :


dy/dx= 1/x+y+1

1
Expert's answer
2022-01-30T14:32:20-0500

From the given question, we have;

"\displaystyle
\frac{dy}{dx}=\frac{1}{x+y+1}"

"\displaystyle
\Rightarrow \frac{dx}{dy}=\frac{x+y+1}{1}=x+y+1\\
\Rightarrow \frac{dx}{dy}-x=(y+1)\qquad\cdots\cdots\cdots\cdots(1)"

which is a linear differential equation with x as dependent variable and y as independent variable.


Comparing (1) with the general form "\displaystyle
\frac{dx}{dy}+Px=Q", which is a linear differential equation with x as dependent variable yields; "\displaystyle
P=-1,\ \text{and } Q=y+1."


Now, the integrating factor (I.F.) "\displaystyle
=e^{\int Pdy}=e^{-\int dy}=e^{-y}"

Multiplying (1) by I.F. yields;

"\displaystyle
\frac{dx}{dy}e^{-y}-xe^{-y}=(y+1)e^{-y}\\
\Rightarrow \frac{d(xe^{-y})}{dy}=(y+1)e^{-y}\\
\Rightarrow e^{-y}x=\int(y+1)e^{-y}\ dy=−ye^{−y}−2e^{−y}+k", where k is an arbitrary constant.

"\displaystyle
\Rightarrow e^{-y}x=-e^{-y}(y+2)+k\\
\Rightarrow x=ke^y-(y+2)"